# Constraining accretion-ejection parameters of a newly discovered Galactic transient Swift J151857.0-572147 using IXPE and NuSTAR observations

### v2

@@ Abstract
We study the spectro-polarimetric properties of a newly discovered black hole X-ray binary Swift J151857.0-572147 jointly using IXPE and NuSTAR observations during March 2024. The analysis of IXPE data reports mode-independent polarization degree (PD) and polarization angle (PA) , while the model-dependent analysis gives PD and PA . The joint spectral analysis of the broadband data and NuSTAR analysis in isolation constrain the mass of the central black hole between and a moderate spin parameter of with disk inclination . The power-law photon index and cutoff energy are and keV, suggesting a transition to the soft spectral state of the source during the observation period. Additionally, the best-fitted halo mass accretion is less compared to the disk mass accretion rate, and a relatively lower corona size of 6-9 rg indicates the same spectral state. The mass outflow rate was also low (% ) during these epochs. The hydrogen column density obtained from the fit is relatively high cm-2.
@@ 1 Introduction
Matter accreting onto a black hole (BH) forms an accretion disk consisting mainly of two components, the inner puffed-up region which is known as the so-called corona, and the cold Keplerian disk. The first component upscatters soft photons from the Keplerian disk through inverse Comptonization making them hard (Sunyaev & Titarchuk, 1980; Haardt & Maraschi, 1993; Chakrabarti & Titarchuk, 1995; Done et al., 2007; Mondal & Chakrabarti, 2013, and references therein). The same region also launches jet/outflows (Chakrabarti, 1999; Chattopadhyay & Das, 2007; Mondal & Chakrabarti, 2021) and polarizes the radiation depending on its temperature, optical depth, geometry, and inclination (Connors et al., 1980; Sunyaev & Titarchuk, 1985). Therefore, the polarization properties such as polarization angle (PA) and degree (PD) of any source depend on both corona properties and the seed photon energies. The variation of PD and PA on the optical depth, disk inclination, and outflows was studied in detail both in theory and using simulation (Dovčiak et al., 2008; Li et al., 2009; Schnittman & Krolik, 2010; Begelman & McKee, 1983, and references therein). Therefore, by constraining the corona properties of a system, the PD can be constrained or vice-versa along with the intrinsic parameters of the BH (Krawczynski et al., 2022; Veledina et al., 2023; Kushwaha et al., 2023; Rawat et al., 2023; Poutanen et al., 2023; Jana & Chang, 2024; Mondal et al., 2024), making spectro-polarimetric studies important in recent days.
@@
Black Hole X-ray binaries (BHXRBs) show variabilities in their spectral states during the outburst time (Remillard & McClintock, 2006). In general, they follow a cycle starting from hard state (HS) when spectra are dominated by hard photons from the corona region to soft state (SS) when spectra are dominated by soft photons from the disk through intermediate states (Homan et al., 2001; Belloni et al., 2005; Remillard & McClintock, 2006; Nandi et al., 2012; Debnath et al., 2015; Shui et al., 2021; Jana et al., 2016, and references therein). Sometimes, some of the sources may not follow the cycle for several reasons e.g. lack of supply of matter or lack of viscosity etc. (King, 1998; Chakrabarti et al., 2019; Mondal, 2020; Mondal et al., 2017). Outbursts that do not go to the SS (Tetarenko et al., 2016, and references therein) or pass through all states are called failed-outburst.
@@
The new Galactic transient Swift J151857.0-572147 was first detected by Swift/XRT as a GRB (GRB 20240303A; Kennea et al., 2024a) in Swift Trigger 1218452 111https://gcn.nasa.gov/circulars/35849(GCN 35849). However, its constant bright nature without signs of fading along with its localization in the Galactic plane was later confirmed to be a Galactic transient. From the best localization of the source using XRT immediate on-board localization, the RA, and Dec of the source were found to be RA(J2000) = 15h 18m 57.00s and Dec(J2000) = -57d 21‘ 47.9“ (Kennea et al., 2024b). Follow-up radio observations with the MeerKAT telescope were done (Cowie et al., 2024) at GHz (L-band) with a bandwidth of MHz at a flux density of 10 mJy (Carotenuto & Russell, 2024) on 04 March 2024 for 15 minutes. The inverted radio spectrum (, where ) with the photon index helped designate the nature of the source as consistent with that of an X-ray binary in the hard state, suggesting it may be a neutron star or a black hole. The Australia Telescope Compact Array (ATCA) followed-up the radio observations on 09 March 2024 for minutes at frequencies and GHz simultaneously (Saikia et al., 2024). Their analysis also affirmed that the source was a Galactic black hole. After the ATCA, Swift/XRT performed target of opportunity (ToO) on this source with an exposure of 1000s. It was found that the spectrum is well described by the combination of phenomenological disk black body (DISKBB) and power-law model (POWERLAW) models (Del Santo et al., 2024). These findings also solidified the nature of the source as a black hole. There was INTEGRAL serendipitous detection of the source from 8 to 11 March 2024 (Sguera, 2024). As part of the monitoring program of GRBs in optical and NIR wavelengths, the 60 cm Robotic Eye Mount (REM) telescope monitored the source in both these wavelengths (Baglio et al., 2024).
@@
The Swift/XRT spectral modeling of the source reported a column density () of cm-2 with a power-law photon index () of (see Kennea et al., 2024a). The negative velocity of an HI absorption line imposes the lower limit on the distance to the source to be kpc. While the absence of positive velocity absorption lines towards other sources in the field of the HI absorption for Swift J151857 puts an upper limit on the distance as kpc (Burridge et al., 2024). Due to a large uncertainty in distance estimation, we used an average distance of 10 kpc throughout our analysis for this source. Other information like the mass, inclination, and spin have not been reported about the source yet. From our analysis, we report these intrinsic properties of the source for the first time.
@@
The Imaging X-ray Polarimetry Explorer (IXPE; Weisskopf et al., 2016) has recently observed several black hole X-ray binaries (BHXRBs) and one of them is SwiftJ151857. The NuSTAR (Harrison et al., 2013) has also observed the source simultaneously. In this paper, therefore, we used both observations to carry out broadband spectral and polarimetric studies of the source to understand the spectral states, accretion properties, and intrinsic parameters of the BH. Along with the corona properties, we have also estimated the mass outflow from this system, that connects the disk-corona-outflow with polarization in a single framework.
The paper is organized as follows: in §2 we discuss the observation and data reduction procedure. In §3, we describe the spectro-polarimetric data analysis results, and finally, we draw our conclusions in §4.
@@ 2 Observation and Data Reduction
For NuSTAR analysis, we considered the observations from 17 and 18 March 2024 and have used the standard NUSTARDAS v1.3.1 333https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/ software to extract the data. We generated cleaned event files using the nupipeline task and generated the spectra using nuproducts along with CALDB version 20220331. We considered a region of for the source and for the background using ds9 (Joye & Mandel, 2003). Furthermore, we used the grppha command to group the data with a minimum of 20 counts in each bin. We used the data of IXPE in 2-8 keV and NuSTAR data in keV energy ranges for each respective epoch of observation. For spectral analysis, we used XSPEC444https://heasarc.gsfc.nasa.gov/xanadu/xspec/ (Arnaud, 1996) version 12.12.0.
3 Data Analysis and Results
3.1 Model-independent IXPE Analysis
@@
The results of the model-independent polarimetric analysis are summarized in Table 1. The measured PD and PA in keV are and respectively. We have also performed energy-dependent PD and PA analyses, which are shown in the left and right panels of Fig. 1 respectively. As it can be observed, the polarization measurements decrease with increasing energy and drop below the MDP99 % above 4 keV. This also gives the PD upper limit (99% confidence level, shown in olive dashed line) of and in the 4-6 and 6-8 keV bands. The 1 confidence contours between PA and PD for different energy bands are shown in Fig. 2, obtained using the PCUBE algorithm.
@@ 3.2 Model-dependent IXPE Analysis
For the model-dependent polarization study, we have fitted all Stokes spectra for the epoch I using absorbed POWERLAW+DISKBB model convolved with a constant polarization POLCONST which reads in XSPEC as CONST*TBABS*POLCONST*(POWERLAW+DISKBB). All Stokes spectra for all three DUs in the energy band keV are fitted simultaneously. The above model gives an acceptable fit with . We determine the PA = and the PD (%) = . Both these values are in excellent agreement with the values obtained from PCUBE algorithm. The neutral hydrogen column density is cm-2, obtained from TBABS model. The POWERLAW model photon index () and disk temperature in DISKBB model are and keV respectively. High values of these two parameters may indicate that the source possibly moved to the soft spectral state during this observation epoch. The confidence contours from the model-dependent analysis are shown in Fig. 3. The solid and dashed contours drawn using XSPEC and PCUBE respectively correspond to 1 confidence.
@@ 3.3 Joint IXPE and NuSTAR Analysis
| I+N1 | 1468/1326 | |||||||
| I+N2 | 1177/1145 | |||||||
| N1 | 999/890 | |||||||
| N2 | 729/709 | |||||||

### v3

# The first detection of X-ray polarization in a newly discovered Galactic transient Swift J151857.0-572147

@@ Abstract
We study the spectro-polarimetric properties of a newly discovered black hole X-ray binary Swift J151857.0-572147 jointly using IXPE and NuSTAR observations during March 2024. The analysis of IXPE data reports the first detection of X-ray polarization with degree (PD) and polarization angle (PA) using model-independent approach, while the model-dependent analysis gives PD and PA . The joint spectral analysis of the broadband data and NuSTAR analysis in isolation constrain the mass of the central black hole between and a moderate spin parameter of with disk inclination . The power-law photon index and cutoff energy are and keV, suggesting a transition to the soft spectral state (SS). Additionally, a relatively lower corona size of , a low mass outflow rate (% ), and the best-fitted halo accretion is less compared to the disk accretion rate further confirms the same state. The low PD detected in the SS can be due to repeated scattering inside the dense corona and the dominant emission from the disk, agrees with the low spin and low disk inclination. The hydrogen column density obtained from the fit is relatively high cm-2.
@@ 1 Introduction
Matter accreting onto a black hole (BH) forms an accretion disk consisting mainly of two components, the inner puffed-up region which is known as the so-called corona, and the cold Keplerian disk. The first component upscatters soft photons from the Keplerian disk through inverse Comptonization making them hard (Sunyaev & Titarchuk, 1980; Haardt & Maraschi, 1993; Chakrabarti & Titarchuk, 1995; Done et al., 2007; Mondal & Chakrabarti, 2013, and references therein). The same region also launches jet/outflows (Chakrabarti, 1999; Chattopadhyay & Das, 2007; Mondal & Chakrabarti, 2021) and polarizes the radiation depending on its temperature, optical depth, geometry, and inclination (Connors et al., 1980; Sunyaev & Titarchuk, 1985, and references therein). In addition to the above geometrical configuration and accretion properties, the spin of the BH is a crucial parameter in deciding the amount of polarization as well as its energy dependence. Therefore, the polarization properties such as polarization angle (PA) and degree (PD) of any source depend on several factors including accretion properties and the spin of the BH. The variation of PD and PA on the optical depth, disk inclination, spin of the BH, and outflows was studied in detail both in theory and simulation (Begelman & McKee, 1983; Dovčiak et al., 2008; Li et al., 2009; Schnittman & Krolik, 2010; Krawczynski, 2012; Taverna et al., 2020; Ratheesh et al., 2024, and references therein). Therefore, by constraining the coronal properties of a system, the PD can be constrained or vice-versa along with the intrinsic parameters of the BH, making spectro-polarimetric studies important in recent time.
@@
Black Hole X-ray binaries (BHXRBs) show variabilities in their spectral states during the outburst time (Remillard & McClintock, 2006). In general, they follow a cycle starting from a hard state (HS) when spectra are dominated by hard photons from the corona region to a soft state (SS) when spectra are dominated by soft photons mainly from the disk through intermediate states (Homan et al., 2001; Belloni et al., 2005; Remillard & McClintock, 2006; Nandi et al., 2012; Debnath et al., 2015; Shui et al., 2021; Jana et al., 2016, and references therein). Sometimes, some of the sources may not follow the cycle for several reasons e.g. lack of supply of matter or lack of viscosity etc. (King, 1998; Chakrabarti et al., 2019; Mondal, 2020; Mondal et al., 2017). Outbursts that do not go to the SS (Tetarenko et al., 2016, and references therein) or do not pass through all states are called failed-outburst. In the SS, the disk moves much closer to the BH and the emission from the corona region is negligibly small. As the corona nearly disappears in the SS and its optical depth increases, it can trap more radiation, lowering the probability of hard emission (see Sunyaev & Titarchuk, 1980; Chakrabarti, 1997; Mondal & Chakrabarti, 2013, and references therein). Therefore, the scattering can also affect the amount of polarization (Sunyaev & Titarchuk, 1985).
@@
The new Galactic transient Swift J151857.0-572147 (hereafter SwiftJ151857) was first detected by Swift/XRT as a GRB (GRB 20240303A; Kennea et al., 2024a) in Swift Trigger 1218452 111https://gcn.nasa.gov/circulars/35849(GCN 35849). However, its constant bright nature without signs of fading along with its localization in the Galactic plane was later confirmed to be a Galactic transient. From the best localization of the source using XRT immediate on-board localization, the RA, and Dec of the source were found to be RA(J2000) = 15h 18m 57.00s and Dec(J2000) = -57d 21‘ 47.9“ (Kennea et al., 2024b). Follow-up radio observations with the MeerKAT telescope were done (Cowie et al., 2024) at GHz (L-band) with a bandwidth of MHz at a flux density of 10 mJy (Carotenuto & Russell, 2024) on 04 March 2024 for 15 minutes. The inverted radio spectrum (, where ) with the photon index helped designate the nature of the source as consistent with that of an X-ray binary in the hard state, suggesting it may be a neutron star or a black hole. The Australia Telescope Compact Array (ATCA) followed-up the radio observations on 09 March 2024 for minutes at frequencies and GHz simultaneously (Saikia et al., 2024). Their analysis also affirmed that the source was a Galactic black hole. After the ATCA, Swift/XRT performed target of opportunity (ToO) on this source with an exposure of 1000s. It was found that the spectrum is well described by the combination of phenomenological disk black body (DISKBB) and power-law model (POWERLAW) models (Del Santo et al., 2024). These findings also solidified the nature of the source as a black hole. There was INTEGRAL serendipitous detection of the source from 8 to 11 March 2024 (Sguera, 2024). As part of the monitoring program of GRBs in optical and NIR wavelengths, the 60 cm Robotic Eye Mount (REM) telescope monitored the source in both these wavelengths (Baglio et al., 2024).
@@
The broadband X-ray data of the source from Insight-HXMT (Zhang et al., 2020) was analyzed by Chatterjee et al. (2024) in the intermediate spectral state. Authors detected the presence of type-C quasi-periodic oscillations (QPOs) during the outburst period, which disappeared in the SS while The Imaging X-ray Polarimetry Explorer (IXPE; Weisskopf et al., 2016) observed the source. The Swift/XRT spectral modeling of the source reported a column density () of cm-2 with a power-law photon index () of (see Kennea et al., 2024a). A similar was estimated by Chatterjee et al. (2024). The negative velocity of an HI absorption line imposes the lower limit on the distance to the source to be kpc. While the absence of positive velocity absorption lines towards other sources in the field of the HI absorption for Swift J151857 puts an upper limit on the distance as kpc (Burridge et al., 2024). Due to a large uncertainty in distance estimation, we used an average distance of 10 kpc throughout our analysis for this source. Other information like the mass, inclination, and spin have not been reported about the source yet. From our analysis, we report these intrinsic properties of the source for the first time.
Recently, IXPE has observed several black hole X-ray binaries (BHXRBs) along with SwiftJ151857 and showcased their X-ray polarization detection. In 4U 1630–47 PD % was observed (Cavero et al., 2023; Kushwaha et al., 2023a; Rawat et al., 2023; Ratheesh et al., 2024), in the HS of Swift J1727.8–1613 the PD measured % (Veledina et al., 2023a; Ingram et al., 2023; Podgorný et al., 2024), while a rapid decrease in PD % was observed by Svoboda et al. (2024a) in the SS. In LMC X-3, % polarization was detected (Svoboda et al., 2024b; Majumder et al., 2024) and in the SS of 4U 1957+115 measured PD was % (Marra et al., 2023; Kushwaha et al., 2023b). A 4% PD was detected in Cyg X-1 in the HS and % in the SS (Krawczynski et al., 2022; Jana & Chang, 2024). Very recently, Veledina et al. (2023b) detected a high polarization in Cyg X-3 using IXPE observation. These studies not only detected polarization but also showed the variation in PD in different spectral states, as expected from the simulations. Therefore, there has been renewed interest in studying BHXRBs to further understand their accretion geometry, emission mechanisms, and intrinsic parameters of the central BH. The SwiftJ151857 source was observed by NuSTAR (Harrison et al., 2013) simultaneously with IXPE.
@@
Motivated by recent X-ray polarization detections, which have been used as a powerful tool to understand several emission and accretion properties of BHXRBs, in this paper, we performed spectro-polarimetric studies of SwiftJ151857. We used both IXPE and NuSTAR observations of the source to carry out broadband spectral and polarimetric studies to understand the origin of X-ray polarization, its properties, spectral states, accretion geometry, and intrinsic parameters of the BH. Along with the corona properties, we have also estimated the mass outflow from this system, that connects the disk-corona-outflow with polarization in a single framework. The paper is organized as follows: in §2 we discuss the observation and data reduction procedure. In §3, we discuss the spectro-polarimetric data analysis results, and the origin of the detected polarization, and finally, we draw our conclusions in §4.
@@ 2 Observation and Data Reduction
For NuSTAR analysis, we considered the observations from 17 and 18 March 2024 and have used the standard NUSTARDAS v1.3.1 333https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/ software to extract the data. We generated cleaned event files using the nupipeline task and generated the spectra using nuproducts along with CALDB version 20220331. We considered a region of for the source and for the background using ds9 (Joye & Mandel, 2003). Furthermore, we used the grppha command to group the data with a minimum of 200 counts in each bin. We used the data of IXPE in 2-8 keV and NuSTAR data in keV energy ranges for each respective epoch of observation. For spectral analysis, we used XSPEC444https://heasarc.gsfc.nasa.gov/xanadu/xspec/ (Arnaud, 1996) version 12.12.0. The statistic is used for the goodness of the fit throughout the analysis. The best-fitted spectral plots are further rebinned for visual clarity, it neither affects the model parameter values nor the goodness of the fit.
@@
3 Results and Discussions
3.1 Model-independent IXPE Analysis
The results of the model-independent polarimetric analysis are summarized in Table 1. The measured PD and PA in keV are and respectively. We have also performed energy-dependent PD and PA analyses, which are shown in the left and right panels of Fig. 1 respectively. As it can be observed, the polarization measurements decrease with increasing energy and drop below the MDP99 % above 4 keV. This also gives the PD upper limit (99% confidence level, shown in olive dashed line) of and in the 4-6 and 6-8 keV bands. The 1 confidence contours between PA and PD for different energy bands are shown in Fig. 2, obtained using the PCUBE algorithm. The errors obtained through the PCUBE algorithm do not include systematic (Kislat et al., 2015).
Our detected low PD for the source SwiftJ151857 in the SS is in agreement with the measurements of PD from other BHXRBs in the same state, e.g., % in Swift J1727.8–1613 (Svoboda et al., 2024a) and % in 4U 1957+115 (Marra et al., 2023). Additionally, an energy-dependent PD variation shows a decreasing trend, where the first two energy bands are above or within the MDP, while the higher energy band measurements are below the MDP. Such a decreasing profile in PD has been obtained from theory as well as from numerical simulations and that can be due to the following reasons. The spectra from accretion disks around BHs consist of mainly two types of photons, the low-energy photons that are emitted from the outer region of the disk, where the temperature is lower and the effect of gravity is less important, whereas high-energy emissions are coming from the inner hot region of the disk, where the photon polarization is mostly affected by general relativistic effects. It was shown from numerical simulations that for higher optical depths (see e.g., Dovčiak et al., 2008; Krawczynski, 2012; Taverna et al., 2020; Marra et al., 2023) which is the case in the SS, direct radiation turns out to be polarized perpendicularly to the disk symmetry axis at lower energies, while it slowly decreases at higher energies ( keV) under the effect of the polarization plane rotation. Such turnover energy is decided depending on the spin of the BH and the disk inclination. For higher spinning or inclination systems, turnover occurs at much closer energy and both PA and PD start increasing. However, our source has low inclination and spin (see Sec. 3.4), therefore, the decreasing PA and PD profiles and turnover around 5 keV are well in agreement with the theory and simulation.
Most of the IXPE observations for other BHXRBs showed nearly a constant PA. However, we found a decrease in PA with energy, which is similar to what was observed in 4U 1957+115 (Marra et al., 2023; Kushwaha et al., 2023b). Such a change in PA could be an indication of rotation of PA between 2-8 keV due to the effects of depolarization. Another possibility can be due to the orientation of jets. However, at present, the jet position angle is not well constrained for this source to compare with the detected PA, requiring a further jet study.
@@ 3.2 Model-dependent IXPE Analysis
For the model-dependent polarization study, we have fitted all Stokes spectra for the epoch I using absorbed POWERLAW+DISKBB model convolved with a constant polarization POLCONST which reads in XSPEC as CONST*TBABS*POLCONST*(POWERLAW+DISKBB). All Stokes spectra for all three DUs in the energy band keV are fitted simultaneously. The above model gives an acceptable fit with . We determine the PA = and the PD (%) = . Both these values are in excellent agreement with the values obtained from PCUBE algorithm. The neutral hydrogen column density is cm-2, obtained from TBABS model. The POWERLAW model photon index () and disk temperature in DISKBB model are and keV respectively. High values of these two parameters may indicate that the source possibly moved to the soft spectral state during this observation epoch. The source was very recently studied in a follow-up work by Chatterjee et al. (2024) using Insight-HXMT data and observed the disappearance of QPOs in the SS during the epochs that coincide with IXPE observations, thereby further confirms our finding. The confidence contours from the model-dependent analysis are shown in Fig. 3. The solid and dashed contours drawn using XSPEC and PCUBE respectively correspond to 1 confidence.
Moreover, as the source moved to the SS, the inner disk moved much closer to the BH, and the corona became highly shrunk, 6-9 (, see sec 3.5) in comparison to HS in other BHXRBs (few 100 ; see Mondal et al., 2014; Debnath et al., 2015, and references therein). In the SS, it is widely believed that the accretion disk extends down to the innermost stable circular orbit (ISCO; Reynolds, 2019, for a review) and the spectrum is dominated by photons emitted from the very inner region, where the strong gravity can significantly alter the observed polarization properties (Krawczynski, 2012; Taverna et al., 2020, and references therein). In this regime, the general relativistic effects cause the photon polarization vectors to undergo rotation, resulting in a net depolarization of the emission at infinity and an overall rotation of the PA. These effects are expected to be stronger for the high-energy photons emitted closer to the central BH, introducing the observed energy dependence of PD for this source in the SS. Numerical simulation performed in the literature (Schnittman & Krolik, 2010; Krawczynski, 2012; Taverna et al., 2020; Marra et al., 2023), showed a similar decreasing profile of PD and PA with energy and that also explains the case of low-spinning BH, which is valid for our source. Therefore, we do expect such a profile in PD for our source, if possible to other sources in the future, during IXPE era. However, for a rapidly spinning BH with high inclination, an opposite PD profile is expected that agrees with the observed PD profile in e.g. 4U 1957+115 in the SS.
We note that polarization measurement changes with spectral states, as we mentioned earlier. Two sources showed a significant change in PD in different spectral states, one is Cyg X-1 (Krawczynski et al., 2022; Jana & Chang, 2024) and the other Swift J1727.8-1613 (Svoboda et al., 2024a). These sources also showed significant detection of polarization in the SS, however, PD in the SS was much less compared to the HS. The number of scattering in the SS increases as photons get trapped by the scattering medium, thereby, polarization may decrease. Podgorný et al. (2023) detected % polarization in the SS of LMC X-1 and preferred a slab geometry of the corona. Our detection of PD in the SS is in accord with the above findings and further supports that the changes in spectral states are possibly followed by the changes in polarization properties. The possible origin of such low polarization favors the scenario in which the X-ray polarization is driven by the geometry of the accretion flow in the innermost region and an optically thick corona. Also, the polarization of the disk is perpendicular to that of the corona. As the disk emission dominates in the SS, the PD could decrease as was explained in Krawczynski et al. (2022).
@@ 3.3 Joint IXPE and NuSTAR Analysis
| I+N1 | 1065/940 | |||||||
| I+N2 | 872/825 | |||||||
| N1 | 596/504 | |||||||
| N2 | 421/389 | |||||||
3.4 KERRBB Model
In Fig. 4, we show the model-fitted spectra of Swift J151857 for the joint IXPE and NuSTAR observations. The left and right panels show the spectral fits for the epochs I+N1 and I+N2. The lower panels show the residuals. As we did not have the mass, spin parameters of the BH, and the disk inclination beforehand, we left them free during model fitting, keeping the distance to the source fixed at 10 kpc. However, parameter values may change if the source distance changes significantly. Both data fits returned consistent values of all three parameters in both epochs. The mass () and spin () parameters estimated from the fit vary between and respectively. It possibly indicates that the BH is moderately rotating. The disk inclination () and the mass accretion rate obtained from the fit vary between and ( gm s-1. The low cutoff energy () of keV with a photon index () of indicates a transition to the soft spectral state of the source during the observation epoch after ejecting a high radio flare in the first week of March. The value of the source was high cm-2, which can be due to strong outflows from the outer disk or due to the presence of some blobs along the line of sight that can block the central radiation (see Neilsen & Homan, 2012; Mondal & Jithesh, 2023), requires further study of the outflows from the outer disk. We observed a broad Fe K width which varies between keV. It is expected as the source moved to the SS, the disk also moved closer to the BH, therefore, relativistic effects can broaden the line (see e.g. Fabian et al., 1989; Iwasawa et al., 1996; Brenneman & Reynolds, 2006; Mondal et al., 2016, and references therein). How close has the disk moved inward toward the BH? What is the corona geometry (radius and height)? This has been studied in the next section.